### v2

# Strategizing against Q-learners: A Control-theoretical Approach

Abstract
In this paper, we explore the susceptibility of the independent Q-learning algorithms (a classical and widely used multi-agent reinforcement learning method) to strategic manipulation of sophisticated opponents in normal-form games played repeatedly. We quantify how much strategically sophisticated agents can exploit naive Q-learners if they know the opponents’ Q-learning algorithm. To this end, we formulate the strategic actors’ interactions as a stochastic game (whose state encompasses Q-function estimates of the Q-learners) as if the Q-learning algorithms are the underlying dynamical system. We also present a quantization-based approximation scheme to tackle the continuum state space and analyze its performance for two competing strategic actors and a single strategic actor both analytically and numerically.
I INTRODUCTION
The widespread adoption of (reinforcement) learning algorithms in multi-agent systems has significantly enhanced autonomous systems to tackle complex tasks through learning from interactions within a shared environment. However, strategically sophisticated actors can exploit such algorithms to perform sub-optimally [1, 2, 3, 4, 5]. A critical question is how much a strategically sophisticated agent can manipulate the opponent’s decisions for more payoff in a game if the agent is aware of the opponent’s learning dynamic. This strategic act can also have a positive impact on the opponent depending on how aligned their objectives are.
To address the challenge, we approach the problem from a control-theoretical perspective. We focus on the repeated play of general-sum normal-form games played by two types of agents: strategically advanced (A-type) and naive (N-type). Each N-type naively follows the widely-used independent Q-learning (IQL) algorithm as if no other agents exist in the environment [6, 7]. On the other hand, an A-type is a strategically sophisticated agent with complete model knowledge including underlying game structure, N-types’ algorithms and observations of all actions. We consider A-types with these advanced capabilities to establish a benchmark for quantifying the performance of other strategic actors with limited capabilities.
We show that A-types can leverage their model knowledge to control/drive N-types’ algorithms (as if the algorithms are some dynamical system) to maximize the discounted sum of the payoffs collected over an infinite horizon. We can model the interactions among A-types as stochastic games (SG)–a generalization of Markov decision processes (MDPs) to non-cooperative multi-agent environments [8]. However, the SG has a continuous state space. To address this issue, as an example , we present a quantization-based approximation scheme reducing the problem to a finite SG that can be solved via standard dynamic programming methods when there are two competing A-types or a single A-type. We quantify the approximation level of this quantization scheme and examine its performance numerically.
Related Works. Strategizing against naive learning algorithms has been studied for no-regret learning and fictitious play dynamics [3, 5, 4]. In [3], the authors show that sophisticated agents can secure the Stackelberg equilibrium value against no-regret learners. In [5], the authors study strategizing against a fictitious player in two-agent games with two actions. In [4], the authors prove that sophisticated agents with the knowledge of game matrix can attain better payoffs than the one at Nash equilibrium by solving a linear program for each action of the fictitious player to obtain her own mixed strategy, then play a pure action trajectory to satisfy the desired mixed strategy probabilities. On the other hand, here, we focus on strategizing against the widely-studied IQLs, e.g., see [9, 10, 11, 12, 7], and bringing the problem to the control-theoretical framework.
Learning algorithms relying on feedback from the environment (such as Q-learning) can be vulnerable to exploitation through the manipulation of the feedback by some adversaries. In a broader sense, our approach relates to the literature on reward poisoning [1, 2, 13, 14]. The main distinction between our work and this body of literature is the presence of an underlying game structure, which constraints the players’ impact on each other. Furthermore, while the general setting in reward poisoning is highly competitive (e.g., zero-sum), we address arbitrary general-sum games. The most relevant studies to our work are those by [1, 2], where the authors investigate the impact of manipulating Q-learner algorithms through the falsification of cost signals and its effect on the algorithm’s convergence. Correspondingly, robust variants of the Q-learning against such attacks have also been studied extensively [15, 16, 17].
Motivating Example. Recently, IQLs have been observed to lead to tacit collusion that can undermine the competitive nature of the markets [11, 12, 18, 19]. For example, in [19], the authors study the collusive behavior of IQLs in the widely-studied prisoner’s dilemma game (where agents have two actions: ‘cooperate’ and ‘deflect’). They observe that Q-learners can learn to collude in cooperation even though ‘cooperate’ is always an irrational choice as ‘deflect’ dominates ‘cooperate’ strategy. New regulations are needed to prevent such collusive behavior [11]. Here, we examine the susceptibility of IQLs to strategic manipulation in prisoners’ dilemma numerically. Such vulnerabilities can incentivize businesses not to use IQLs with the hope of tacit collusion as their naive approach might be exploited by the other business with strategically sophisticated algorithms.
Contributions. Our main contributions include:
  • To the best of our knowledge, our work is the first to model the vulnerability of IQLs to strategic manipulations by advanced players as an SG as if the IQLs are the underlying dynamic environment.
  • The continuum state space in the SGs poses a technical challenge. However, we demonstrate the Lipschitz continuity of the value functions for the SGs based on the specifics of the IQL updates. This characterization shows that the SG formulated is relatively well-behaved as the approximation methods, such as the value-based ones, can perform effectively.
Organization. The paper is organized as follows: We formulate the strategic actors’ interactions as an SG in Section II. We present the solution approach in Section III. We provide illustrative examples and conclude the paper, resp., in Sections IV and V. Three appendices include the proofs of technical lemmas.
Figure 1: Illustration of the interaction between N-types and A-types across the repeated play of the normal-form game $\mathcal{G}$. A-types compete against each other while considering N-types’ q-function estimates as an underlying state of an SG. A-types can use dynamic programming to solve the SG and use trackers to track N-types’ q-function estimates by leveraging the complete model knowledge.
II STRATEGIC ACTORS AGAINST Q-LEARNERS
Consider a normal-form game $\mathcal{G}$ characterized by the tuple $\langle\{A^{j}\}_{j\in I^{A}},\{B^{i}\}_{i\in I^{N}},\{u^{j}\}_{j\in I^{A}},\{\overline{u}^{i}\}_{i\in I^{N}}\rangle$, where $I^{A}$ and $I^{N}$ denote, resp., the index set of A-type and N-type agents.
  • N-types are naive players following the IQL algorithm based on the local information (i.e., payoff received) as if there are no other agents.
  • A-types are advanced players with the complete knowledge of the underlying game and the employed IQLs. They can also access to the joint actions of all players.
Let $A^{j}$ and $B^{i}$ denote the A-type $j$’s and N-type $i$’s finite action sets. The joint action spaces of A-types and N-types are denoted by $A\coloneqq\times_{j\in I^{A}}A^{j}$ and $B\coloneqq\times_{i\in I^{N}}B^{i}$, respectively. The payoff functions are $u^{j}:A\times B\rightarrow\mathbb{R}$ and $\overline{u}^{i}:A\times B\rightarrow\mathbb{R}$. A-types and N-types are repeatedly playing $\mathcal{G}$ by taking actions simultaneously over stages $k=0,1,\ldots$ as follows:
N-type: Let $q_{k}^{i}:B^{i}\rightarrow\mathbb{R}$ denote N-type $i$’s q-function estimate at stage $k$. Then, the agent takes action according to the softmax function:
$\sigma(q_{k}^{i})(b^{i})=\frac{\exp(q_{k}^{i}(b^{i})/\tau)}{\displaystyle\sum_{\tilde{b}^{i}\in B^{i}}\exp(q_{k}^{i}(\tilde{b}^{i})/\tau)}\quad\forall b^{i}\in B^{i},$ (1)
for some $\tau>0$ controlling the level of exploration and $\sigma(q_{k}^{i})(b^{i})\in(0,1]$ denotes the probability that they take action $b^{i}\in B^{i}$. Let $a_{k}=(a_{k}^{j})_{j\in I^{A}}$ and $b_{k}=(b_{k}^{i})_{i\in I^{N}}$ denote the joint actions of A-types and N-types at stage $k$, respectively. Given the payoff $\overline{u}_{k}^{i}=\overline{u}^{i}(a_{k},b_{k})$, N-type $i$ updates the q-function according to
$q_{k+1}^{i}(b^{i})=\begin{cases}q_{k}^{i}(b^{i})+\alpha(\overline{u}_{k}^{i}-q_{k}^{i}(b^{i}))&\;\text{if }b^{i}=b_{k}^{i}\\ q_{k}^{i}(b^{i})&\;\text{o.w}\end{cases},$ (2)
with some step size $\alpha\in(0,1)$ and the initial vector $q_{0}^{i}(b^{i})=0$ for all $b^{i}$. [^1] We interchangeably view q-function as a function $q^{i}:B^{i}\rightarrow\mathbb{R}$ and a vector $q^{i}\in\mathbb{R}^{|B^{i}|}$ since $B^{i}$ is a finite set. In other words, N-types act myopically as if the other players play according to some stationary strategy across the repeated play of the underlying game $\mathcal{G}$.
A-type: The advanced players can track N-types’ q-function estimates by knowing the deployed IQLs, as described in (2), and observing the actions of all agents. Therefore, A-types do not act myopically as they are aware that the other players do not necessarily play according to some stationary strategy. Instead, they strategically aim to maximize the discounted sum of the payoffs they can collect across the repeated play of the underlying game $\mathcal{G}$, i.e.,
$\max_{\{a_{k}^{i}\}_{k=1}^{\infty}}\mathrm{E}\left[\sum_{k=0}^{\infty}\gamma^{k}u^{i}\left(a_{k},b_{k}\right)\right],$ (3)
for some discount factor $\gamma\in(0,1)$.
A-type $i$ takes actions to achieve (3) based on the history $\{a_{0},b_{0},\ldots,a_{k-1},b_{k-1}\}$ of the game and the knowledge that $q_{k}^{i}$’s evolve according to (2). However, this is not a well-defined optimization problem that can be solved introspectively due to its dependence on the other players’ play.
Although the information set $\{a_{0},b_{0},\ldots,a_{k-1},b_{k-1}\}$ grows unboundedly in time, N-types’ actions depend only on their q-function estimates. In other words, conditioned on the current q-function estimates and the current action profile of A-types, the next q-function estimates, and therefore, N-types’ play is independent of the game history. A-types can leverage this Markov property in their objective (3) by viewing the q-function estimates as some underlying controlled Markov process. Therefore, we can model A-types’ interactions with each other as the following stochastic game whose state is N-types’ q-function estimates.
Problem 1 (Stochastic Game Formulation) Consider an $|I^{A}|$-agent SG characterized by the tuple $\mathcal{Z}=\langle Z,(A^{j},r^{j})_{j\in I^{A}},\gamma,p\rangle$. The state space $Z\coloneqq\prod_{i\in I^{N}}Q^{i}$ is a compact set, where $Q^{i}\subset\mathbb{R}^{|B^{i}|}$ corresponds to all possible q-function estimates of N-type $i\in I^{N}$. The reward function $r^{j}:Z\times A\rightarrow\mathbb{R}$ is given by
$r^{j}(z,a)=\mathrm{E}_{b\sim\overline{\pi}(z)}[r^{j}(a,b)]\quad\forall(z,a)\in Z\times A,$ (4)
where
$\overline{\pi}(\{q^{i}\}_{i\in I^{N}})(\{b^{i}\}_{i\in I^{N}})\coloneqq\prod_{i\in I^{N}}\sigma(q^{i})(b^{i})$ (5)
for all $q^{i}\in Q^{i}$, $b^{i}\in B^{i}$ and $i\in I^{N}$. [^2] By concatenation of $\{q^{i}\}_{i\in I^{N}}$ we can construct state $z$. Since $b$ is the action profile, $\overline{\pi}(\{q^{i}\}_{i\in I^{N}})(\{b^{i}\}_{i\in I^{N}})$ is equivalent to $\overline{\pi}(z)(b)$. The transition kernel $p(\cdot\mid\cdot)$ determines the evolution of transformed states according to the update rule (2).
Recall that A-types know N-types’ IQLs (2) and the underlying game model $\mathcal{G}$ and can observe the actions $(a_{k},b_{k})$ of all agents. Therefore, they can track the q-function estimates perfectly although they do not have direct access to them. For example, A-types can deploy a q-tracker $\widehat{q}_{k}^{i}$ specific to each N-type $i$ evolving according to
$\widehat{q}_{k+1}^{i}(b^{i}_{k})=\widehat{q}_{k}^{i}(b^{i}_{k})+\alpha(\overline{u}^{i}(a_{k},b_{k})-\widehat{q}_{k}^{i}(b^{i}_{k}))$ (6a)
$\widehat{q}_{k+1}^{i}(b^{i})=\widehat{q}_{k}^{i}(b^{i})\quad\forall b^{i}\neq b_{k}^{i}$ (6b)
with $\widehat{q}_{0}^{i}=q_{0}^{i}$. Therefore, we have $\widehat{q}_{k}^{i}\equiv q_{k}^{i}$ for all $k$ and A-types can construct the state $z_{k}=(\widehat{q}_{k}^{i})_{i\in I^{N}}$ of the underlying SG $\mathcal{Z}$.
Let each A-type $j$ follow Markov stationary strategy $\pi^{j}:Z\rightarrow\Delta(A^{j})$ in the SG $\mathcal{Z}$. Given the strategy profile $\pi=(\pi^{j})_{j\in I^{A}}$, A-type $j$’s utility (3) can be rewritten as
$U^{j}(\pi^{j},\pi^{-j}):=\mathrm{E}_{a_{k}\sim\pi(z_{k})}\left[\sum_{k=0}^{\infty}\gamma^{k}r^{j}(z_{k},a_{k})\right].$ (7)
for the SG $\mathcal{Z}$. As we pointed out earlier, finding the best strategy in games is generally not a well-defined optimization problem. Therefore, we focus on the important special case of two competing A-types. Let $I^{A}=\{j_{1},j_{2}\}$ and A-types have opposite payoffs $u^{j_{1}}+u^{j_{2}}\equiv 0$ such that $U^{j_{1}}+U^{j_{2}}\equiv 0$.
Algorithm 1 A-type $j$
  input: $\pi^{j}$ for the SG $\mathcal{Z}$ and the initials $\{{q}_{0}^{i}\}_{i\in I^{N}}$
  for each stage $k=0,1,\ldots$ do
     construct $z_{k}=\{{q}_{k}^{i}\}_{i\in I^{N}}$
     take action $a_{k}^{j}\sim\pi^{j}(z_{k})$ simultaneously with the others
     receive payoff $u^{j}(a_{k},b_{k})$ (corresponding to $r^{j}(z_{k},a_{k})$ in expectation)
     observe $(a_{k},b_{k})$
     update the q-trackers according to (6)
  end for
Algorithm 2 N-type $i$
  input: the initial $q_{0}^{i}$
  for each stage $k=0,1,\ldots$ do
     take action $b_{k}^{i}\sim\sigma(q_{k}^{i})$ simultaneously with the others
     receive payoff $\overline{u}_{k}^{i}=\overline{u}^{i}\left(a_{k},b_{k}\right)$
     update the q-function estimate according to (2)
  end for
Definition 1 (Markov Stationary Equilibrium) We say that a Markov stationary profile $\pi=(\pi^{j}:Z\rightarrow\Delta(A^{j}))_{j\in I^{A}}$ is Markov stationary equilibrium provided that [^3] The probability simplex over a finite set $A$ is denoted by $\Delta(A)$.
$U^{j}(\pi^{j},\pi^{-j})\geq U^{j}(\tilde{\pi}^{j},\pi^{-j})\quad\forall\tilde{\pi}^{j},$ (8)
where $\pi^{-j}=\{\pi^{\ell}\}_{\ell\neq j}$ [8].
Proposition 1. There always exists Markov stationary equilibrium for two-agent zero-sum SGs $\mathcal{Z}=\langle Z,(A^{j},r^{j})_{j\in I^{A}},\gamma,\overline{p}\rangle$.
Proof:
The proof follows from [20, Theorem 6.2.12] based on the observation that the minimax function is non-expansive like the maximum function, and the set of all possible q-function estimates, $Z$, is a Polish space as a compact subset of $\mathbb{R}^{|B|}$ and the set of actions, $A$, is finite and state-invariant. ∎
Remark 1. If there is only a single A-type, then the SG $\mathcal{Z}$ reduces to an MDP, for which there also exists a Markov stationary solution.
Given the Markov stationary strategy $\pi^{j}$ for the SG $\mathcal{Z}$, A-type $j$ interacts with other agents as illustrated in Fig. 1. Detailed descriptions of the A-types’ and N-types’ interactions across the repeated play of the underlying game $\mathcal{G}$ are tabulated, resp., in Algorithms 1 and 2.
III SOLVING THE STOCHASTIC GAME
A critical challenge for finding the best strategy in the two-agent zero-sum SG $\mathcal{Z}$ is the continuum state space $Z$. Based on the specifics of the IQLs, we can show that the SG $\mathcal{Z}$ is relatively well-behaved for effective approximations. Particularly, let $v_{\kappa}^{j}:{Z}\rightarrow\mathbb{R}$ denote the values of the states in $Z$ at the initial stage of a $(\kappa+1)$-length horizon for the best decision rules in $\mathcal{Z}$. More explicitly, for each $z\in Z$, we have
$v_{\kappa}^{j_{1}}({z})=\max_{\mu^{j_{1}}}\min_{\mu^{j_{2}}}\mathrm{E}_{(a^{j_{1}},a^{j_{2}})\sim(\mu^{j_{1}},\mu^{j_{2}})}[Q_{\kappa}^{j_{1}}(z,a^{j_{1}},a^{j_{2}})]$ (9a)
$v_{\kappa}^{j_{2}}({z})=\max_{\mu^{j_{2}}}\min_{\mu^{j_{1}}}\mathrm{E}_{(a^{j_{1}},a^{j_{2}})\sim(\mu^{j_{1}},\mu^{j_{2}})}[Q_{\kappa}^{j_{2}}(z,a^{j_{1}},a^{j_{2}})],$ (9b)
where
$Q^{j}_{\kappa}(z,a)=r^{j}({z},a)+\gamma\int_{{Z}}v_{\kappa-1}^{j}({\tilde{z}})p(d{\tilde{z}}\mid{z},a)$ (10)
for all $\kappa>0$, $j\in I^{A}$ and $(z,a)\in Z\times A$. The initial $Q_{0}^{j}({z},a)=r^{j}({z},a)$ for all $(z,a)$.
Proposition 2. The minimax value function $v_{\kappa}^{j}(z)$ for each $j\in I^{A}$ is bounded by
$M_{\kappa}:=\frac{1-\gamma^{\kappa+1}}{1-\gamma}\max_{(\tilde{j},a,b)}|u^{\tilde{j}}(a,b)|$ (11)
for all $z\in Z$ and $\kappa\geq 0$. The value function is also $L_{\kappa}$-Lipschitz with respect to $\|\cdot\|_{1}$, where
$L_{\kappa}=\gamma L_{\kappa-1}+L_{0}\frac{1-\gamma^{\kappa+1}}{1-\gamma}$ (12)
for all $\kappa>0$ and $L_{0}\coloneq\frac{1}{\tau}\sqrt{|B|}\max_{\tilde{j},a,b}|u^{\tilde{j}}(a,b)|$.
Proof:
The boundedness follows from the definitions (4) and (9) since we can bound $v_{\kappa}^{j}(\cdot)$ by
$|v_{\kappa}^{j}({z})|$ $\leq\sum_{k=0}^{\kappa}\gamma^{k}\max_{({z},a)}|r^{j}({z},a)|$
$\leq\sum_{k=0}^{\kappa}\gamma^{k}\max_{(\tilde{j},a,b)}|u^{\tilde{j}}(a,b)|$
$\leq\frac{1-\gamma^{\kappa+1}}{1-\gamma}\max_{(\tilde{j},a,b)}|u^{\tilde{j}}(a,b)|=:M_{\kappa}.$ (13)
Next, we first present the following lemma playing an important role in showing the Lipschitz continuity of the value function.
Lemma 1. The joint strategy $\overline{\pi}(\cdot)$ of N-types, as described in (5), satisfies
$\|\overline{\pi}(z)-\overline{\pi}(\overline{z})\|_{1}\leq\frac{\sqrt{|B|}}{\tau}\cdot\|z-\overline{z}\|_{1}$ (14)
for all $z,\overline{z}\in Z$.
The second part of the proof follows from induction. Firstly, for $\kappa=0$, the non-expansiveness of the minimax function leads to
$|v_{0}^{j}({z})-$ $v_{0}^{j}({\overline{{z}}})|$
$\leq\max_{\mu^{j}}\min_{\mu^{-j}}\;\mathrm{E}[|r^{j}({z},a^{j},a^{-j})-r^{j}({\overline{z}},a^{j},a^{-j})|],$ (15)
where the expectation is taken with respect to $a^{j}\sim\mu^{j}$ and $a^{-j}\sim\mu^{-j}$. Then, we can use the following lemma showing the Lipschitz continuity of the reward functions $r^{j}(\cdot)$’s.
Lemma 2. The reward function $r^{j}({z},a)$ is Lipschitz continuous in $z\in Z$. For each $a\in A$, we have
$|r^{j}({z},a)-r^{j}({\overline{z}},a)|\leq L_{0}\cdot\|{\color[rgb]{0,0,0}z-\overline{z}}\color[rgb]{0,0,0}\|_{1}$ (16)
for all ${z,\overline{z}\in Z}$, where $L_{0}$ is as described in Proposition III.
By (III), Lemma III yields that
$|v_{0}^{j}(z)-v_{0}^{j}(\overline{z})|$ $\leq\max_{\mu^{j}}\min_{\mu^{-j}}\mathrm{E}[|r^{j}({z},a)-r^{j}({\overline{z}},a)|]$
$\leq L_{0}\cdot\|z-\overline{z}\|_{1}.$ (17)
Next, we focus on $\kappa>0$ and assume that $v_{\kappa-1}^{j}(\cdot)$ is $L_{\kappa-1}$-Lipschitz with respect to $\|\cdot\|_{1}$ to show that $v_{\kappa}^{j}(\cdot)$ is $L_{\kappa}$-Lipschitz. The non-expansiveness of the minimax function and the triangle inequality yield that
$|v_{\kappa}^{j}(z)-v_{\kappa}^{j}(\overline{z})|\leq\max_{\mu^{j}}\min_{\mu^{-j}}\mathrm{E}[|r^{j}(z,a)-r^{j}(\overline{z},a)|]$
$+\gamma\max_{\mu^{j}}\min_{\mu^{-j}}\mathrm{E}\left[\left|\int_{Z}v_{\kappa-1}^{j}(\tilde{z})p(d\tilde{z}|z,a)-\int_{Z}v_{\kappa-1}^{j}(\tilde{z})p(d\tilde{z}|\overline{z},a)\right|\right].$
For the first term on the right-hand side, we can invoke Lemma III. On the other hand, for the second term, we can use the following lemma.
Lemma 3. Let $\eta:Z\rightarrow\mathbb{R}$ be bounded $|\eta(z)|\leq M$ for all $z\in Z$ and $K$-Lipschitz with respect to $\|\cdot\|_{1}$. Then we have
$\left|\int_{Z}\eta(\tilde{z})p(d\tilde{z}|z,a)-\int_{Z}\eta(\tilde{z})p(d\tilde{z}|\overline{z},a)\right|\leq\overline{K}\cdot\|z-\overline{z}\|_{1},$ (18)
where $\overline{K}=K+\frac{\sqrt{|B|}}{\tau}\cdot M$.
Based on the boundedness of the value function (13) and the assumption that $v_{\kappa-1}^{j}$ is $L_{\kappa-1}$-Lipschitz, we obtain
$|v_{\kappa}^{j}(z)-v_{\kappa}^{j}(\overline{z})|$ $\leq\left(L_{0}+\gamma\left(L_{\kappa-1}+\frac{\sqrt{|B|}}{\tau}M_{\kappa-1}\right)\right)\cdot\|z-\overline{z}\|_{1}$
$=L_{\kappa}\cdot\|z-\overline{z}\|_{1}$ (19)
for all $z,\overline{z}\in Z$, where the equality follows from the definitions of $M_{\kappa-1}$, $L_{\kappa}$, and $L_{0}$ in the proposition statement. This completes the proof. ∎
The Lipschitz continuity of the minimax value function is promising to effectively approximate the underlying SG via the value-based approximation methods. Various methods exist for approximating MDPs with large state-action spaces. For instance, state aggregation [21, 22] involves combining similar states with respect to cost and transition probabilities into meta-states. On the other hand, in approximate linear programming [23, 24], preselected basis functions are used to fit the value function of the MDP. Approximate value iteration and approximate policy iteration [25, 26, 27] also use preselected basis functions to determine the value and policy, respectively. Neuro-dynamic programming [26] leveraging the power of approximation structures, such as neural networks, is also an important method to approximate MDPs.
In this paper, as an example proof of concept, we focus on quantization-based approximations, e.g., see [28] for a detailed review, due to their interpretability by transforming $\mathcal{Z}$ to a finite SG, as in the original definition in [8]. To this end, we consider a quantization mapping $\Phi:Z\rightarrow D$ for some finite set $D\subset Z$ such that $Z_{d}:=\{z\in Z:\Phi(z)=d\}$ satisfies $Z_{d}\cap Z_{d^{\prime}}=\varnothing$ for all $d\neq d^{\prime}$ and $\bigcup_{d\in D}Z_{d}=Z$. In the following, we quantify the approximation error induced by such quantization schemes based on Proposition III.
Proposition 3. Given the quantization mapping $\Phi$, assume that there exists $\Delta<0$ such that $\|z-\Phi(z)\|_{1}\leq\Delta$ for all $z\in Z$. Then, for each $j\in I^{A}$, we have
$\|v_{\kappa}^{j}-\widehat{v}_{\kappa}^{j}\|_{\infty}\leq\frac{\Delta\sqrt{|B|}}{\tau(1-\gamma)^{3}}\max_{\tilde{j},a,b}|u^{\tilde{j}}(a,b)|\quad\forall\kappa\geq 0,$ (20)
where $v_{\kappa}^{j}$ is described in (9), and
$\widehat{v}_{\kappa}^{j}(z)=\max_{\mu^{j}}\min_{\mu^{-j}}\;\mathrm{E}_{(a^{j},a^{-j})\sim(\mu^{j},\mu^{-j})}[\widehat{Q}_{k}^{j}(z,a^{j},a^{-j})],$ (21)
where
$\widehat{Q}_{\kappa}^{j}(z,a)=r^{j}(\Phi(z),a)+\gamma\int_{Z}\widehat{v}_{\kappa-1}^{j}(\tilde{z})p(d\tilde{z}\mid\Phi(z),a).$ (22)
Proof:
We first focus on the case $\kappa=0$. By the definitions of $v_{0}^{j}$ and $\widehat{v}_{0}^{j}$, we have
$|v_{0}^{j}(z)-\widehat{v}_{0}^{j}$ $(z)|=\big{|}\max_{\mu^{j}}\min_{\mu^{-j}}\mathrm{E}[r^{j}(z,a^{j},a^{-j})]$
$\hskip 36.135pt-\max_{\mu^{j}}\min_{\mu^{-j}}\mathrm{E}[r^{j}(\Phi(z),a^{j},a^{-j})]\big{|}$
$\leq\max_{\mu^{j}}\min_{\mu^{-j}}\mathrm{E}\left[\left|r^{j}(z,a^{j},a^{-j})-r^{j}(\Phi(z),a^{j},a^{-j})\right|\right]$
$\leq L_{0}\cdot\|z-\Phi(z)\|_{1}\quad\forall z\in Z,$ (23)
where we use the non-expansiveness of the minimax function and Lemma III. Next, we focus on $\kappa>0$. Based on the definitions of $v_{\kappa}^{j}$ and $\widehat{v}_{\kappa}^{j}$, resp., as described in (9) and (21), the non-expansiveness of the minimax function and the triangle inequality yield that
$|v_{\kappa}^{j}(z)-\widehat{v}_{\kappa}^{j}(z)|$ $\leq\max_{\mu^{j}}\min_{\mu^{-j}}\mathrm{E}\left[|r^{j}(z,a^{j},a^{-j})-r^{j}(\Phi(z),a^{j},a^{-j})|\right]$
$+\gamma\max_{\mu^{j}}\min_{\mu^{-j}}\mathrm{E}\Big{[}\Big{|}\int_{Z}v_{\kappa-1}^{j}(\tilde{z})p(d\tilde{z}\mid z,a^{j},a^{-j})$
$\hskip 7.22743pt-\int_{Z}\widehat{v}_{\kappa-1}^{j}(\tilde{z})p(d\tilde{z}\mid\Phi(z),a^{j},a^{-j})\Big{|}\Big{]}.$
We can bound the first term on the right-hand side by using Lemma III. Therefore, we focus on the second term. However, in the second term, not only the integrands $v_{\kappa-1}^{j}$ and $\widehat{v}_{\kappa-1}^{j}$ but also the measures $p(\cdot\mid z,a)$ and $p(\cdot\mid\Phi(z),a)$ are different. To address this issue, we can add and subtract the term $\int_{Z}v_{\kappa-1}^{j}(\tilde{z})p(d\tilde{z}\mid\Phi(z),a)$. Then, by the triangle inequality, we have
$\left|\int_{Z}v_{\kappa-1}^{j}(\tilde{z})p(d\tilde{z}\mid z,a)-\int_{Z}\widehat{v}_{\kappa-1}^{j}(\tilde{z})p(d\tilde{z}\mid\Phi(z),a)\right|$
$\hskip 14.45377pt\leq\left|\int_{Z}v_{\kappa-1}^{j}(\tilde{z})p(d\tilde{z}\mid z,a)-\int_{Z}v_{\kappa-1}^{j}(\tilde{z})p(d\tilde{z}\mid\Phi(z),a)\right|$
$\hskip 28.90755pt+\left|\int_{Z}(v_{\kappa-1}^{j}(\tilde{z})-\widehat{v}_{\kappa-1}^{j}(\tilde{z}))p(d\tilde{z}\mid\Phi(z),a)\right|.$ (24)
To bound the first term on the right-hand side of (24), we can invoke Proposition III and Lemma III such that
$\left|\int_{Z}v_{\kappa-1}^{j}(\tilde{z})p(d\tilde{z}\mid z,a)-\int_{Z}v_{\kappa-1}^{j}(\tilde{z})p(d\tilde{z}\mid\Phi(z),a)\right|$
$\hskip 57.81621pt\leq\left(L_{\kappa-1}+\frac{\sqrt{|B|}}{\tau}M_{\kappa-1}\right)\|z-\Phi(z)\|_{1}$
$\hskip 57.81621pt=\left(L_{\kappa-1}+L_{0}\frac{1-\gamma^{\kappa}}{1-\gamma}\right)\|z-\Phi(z)\|_{1},$ (25)
where the equality follows from the definitions of $L_{0}$ and $M_{\kappa-1}$. On the other hand, we can bound the second term on the right-hand side of (24) by
$\left|\int_{Z}(v_{\kappa-1}^{j}(\tilde{z})-\widehat{v}_{\kappa-1}^{j}(\tilde{z}))p(d\tilde{z}\mid\Phi(z),a)\right|$
$\hskip 108.405pt\leq\|v_{\kappa-1}^{j}-\widehat{v}_{\kappa-1}^{j}\|_{\infty}.$ (26)
Then, by Lemma III and the bounds (25), (26), we obtain
$|{v}_{\kappa}^{j}(z)-\widehat{v}_{\kappa}^{j}$ $(z)|\leq\left(L_{0}+\gamma L_{\kappa-1}+\gamma L_{0}\frac{1-\gamma^{\kappa}}{1-\gamma}\right)\|z-\Phi(z)\|_{1}$
$\hskip 28.90755pt+\gamma\|v_{\kappa-1}^{j}-\widehat{v}_{\kappa-1}^{j}\|_{\infty}$
$=L_{\kappa}\cdot\|z-\Phi(z)\|_{1}+\gamma\|v_{\kappa-1}^{j}-\widehat{v}_{\kappa-1}^{j}\|_{\infty}.$ (27)
Given $\Delta\geq\|z-\Phi(z)\|_{1}$ for all $z$, we have
$\|v_{0}^{j}-\widehat{v}_{0}^{j}\|_{\infty}\leq L_{0}\cdot\Delta$ (28a)
$\|{v}_{\kappa}^{j}-\widehat{v}_{\kappa}^{j}\|_{\infty}\leq L_{\kappa}\cdot\Delta+\gamma\|v_{\kappa-1}^{j}-\widehat{v}_{\kappa-1}^{j}\|_{\infty}\quad\forall\kappa>0$ (28b)
The recursion (28) on $\|v_{\kappa}^{j}-\widehat{v}_{\kappa}^{j}\|_{\infty}$ implies that
$\|v_{\kappa}^{j}-\widehat{v}_{\kappa}^{j}\|_{\infty}\leq\Delta\cdot\sum_{k=0}^{\kappa}\gamma^{\kappa-k}L_{k}.$ (29)
On the other hand, the recursion on $\{L_{\kappa}\}_{\kappa\geq 0}$ implies that
$L_{\kappa}=L_{0}\sum_{\ell=0}^{\kappa}(\ell+1)\gamma^{\ell}.$ (30)
After some algebra, (29) and (30) yield (20). ∎
Figure 2: The evolution of the empirical averages of the action profiles for Q-learner vs Q-learner in the prisoner’s dilemma.
Figure 3: The evolution of the empirical averages of the action profiles for strategic actor vs Q-learner in the prisoner’s dilemma.
Figure 4: The evolution of the empirical averages of the action profiles for Q-learner vs Q-learner in the two-agent four-action zero-sum game.
Figure 5: The evolution of the empirical averages of the action profiles for strategic actor vs Q-learner in the two-agent four-action zero-sum game.
IV ILLUSTRATIVE EXAMPLES
We illustrate the performance of A-types and N-types interacting with each other in diverse set of scenarios: $(i)$ single A-type against single N-type, $(ii)$ single A-type against multiple N-types, and $(iii)$ two competing A-types against a single N-type. We set the parameters of the IQLs deployed by N-types as $\tau=0.01$ and $\alpha=0.05$. We set $\gamma=0.8$ for the utility (3). We compute the Markov stationary strategy of A-types based on the quantization scheme described in Section III. We quantize the entries of the q-functions to $20$ uniform intervals for the scenarios with two N-types, and $100$ uniform intervals for the rest of the experiments. We run $100$ independent trials and present the mean empirical average of the action profiles as in [19]. In all these scenarios, A-types collected more payoff (and therefore, attained higher utility) on average, compared to the cases where they also deploy IQLs.
Single A-type against Single N-type. We focus on the prisoners’ dilemma game due to its importance for tacit collusion. Furthermore, we study the performance of the proposed quantization-based method in randomly generated two-agent zero-sum game where each agent has four actions. In the prisoners’ dilemma, under the strategic act of A-type, the empirical frequency of cooperation decreased from 95% to 60% since A-type pushed the N-type to play “C” when she plays “D”, as illustrated in Fig. 3. Therefore, the A-type’s strategic act exploits N-type’s IQL and increases A-type’s utility (3) from 0.1041 to 0.2051.444The utilities are normalized with $(1-\gamma)$ for $10^{3}$ iterations. Similarly, in the zero-sum game, A-type achieved the utility of 0.5340 when she acted strategically, which is higher than the value 0.4465 when she deployed IQL.
TABLE I: Single A-type vs Multiple N-types
Agent Type Learning Dynamics Average Value
Aligned Obj. NxNxN N-type IQL 0.1593
N-type IQL 0.1593
N-type IQL 0.1593
NxNxA A-type DP 0.2396
N-type IQL 0.2396
N-type IQL 0.2396
Misaligned Obj. NxNxN N-type IQL -0.1470
N-type IQL 0.1470
N-type IQL 0.1470
NxNxA A-type DP -0.0810
N-type IQL 0.0810
N-type IQL 0.0810
TABLE II: Two Competing A-types vs Single N-type
Agent Type Payoffs Learning Dynamics Average Value
NxNxN N-type $\widetilde{U}$ IQL 0.4280
N-type $U$ IQL 0.3640
N-type $-U$ IQL -0.3640
NxAxN N-type $\widetilde{U}$ IQL 0.4875
A-type $U$ DP 0.4995
N-type $-U$ IQL -0.4995
AxAxN N-type $\widetilde{U}$ IQL 0.4854
A-type $U$ Minimax-DP 0.5092
A-type $-U$ Minimax-DP -0.5092
When the number of players are scaled to three, A-type continues to obtain a higher value compared to the cases where she also deploys IQLs.
Single A-type against Multiple N-types. We perform experiments on a single A-type and two N-types for the cases where A-type has aligned and misaligned objectives with N-types and N-types play a potential game given any action of A-type. For the aligned objectives, we observe that when A-type acts strategically by solving the MDP (see Remark 2), the average value of all agents increases, indicating that A-type drives N-types to higher payoffs. For the misaligned objectives, A-type’s payoff is the opposite of the N-types’ potential function. We observe that A-type can exploit N-types for its own benefit while N-types suffer from the A-type’s strategic act. These results, depicted in Table I, quantifies how strategically sophisticated agents can manipulate naive agents in potential games both positively and negatively depending on the alignments of their objectives.
Two competing A-types against Single N-type. We let N-type have an arbitrary payoff matrix, denoted by $\widetilde{U}$, while A-types have opposite objectives, denoted by $U$ and $-U$. In Table II, we first report the average values for the case NxNxN where all agents adopt IQLs. Then, we demonstrate the effect of a strategically sophisticated agent by allowing one to have A-type, i.e., act strategically by solving the MDP via DP. Note that now N-types does not play a potential game given any action of A-type, different from the example above. Depending on the alignment of their objectives with A-type, the strategic act of A-type has positive and negative impact on them. Lastly, we show the impact of competing A-types deploying minimax-DP to compute their minimax-equilibrium strategy for the two-agent zero-sum SG $\mathcal{Z}$. Interestingly, N-type benefits from the competition between A-types when they act strategically. We attribute this to the alignment of N-type’s payoff with A-type achieving higher payoff in AxAxN than NxNxN. Hence, we can observe such intriguing outcomes for competing strategic actors.
V CONCLUSION
We addressed how strategically sophisticated agents can strategize against naive opponents following independent Q-learning algorithms in normal-form games played repeatedly if the strategic actors know the deployed Q-learning algorithms and the underlying game payoffs. We modeled the strategic actors’ interactions as an SG as if the Q-learning algorithms are the underlying dynamic environment. However, this resulted in a SG with a continuum state space. We demonstrated the value function specific to Q-learning updates in the continuum state space is Lipschitz continuous, promising for effective approximations via value-based methods. As an example proof of concept, we focused on quantization-based approximation to reduce the problem to a finite SG that can be solved via minimax value iteration when there are two competing strategic actors. We also analytically quantified the approximation level and numerically examined the quantization scheme’s performance.
This work paves the way for understanding the vulnerabilities of learning algorithms against strategic actors from a control-theoretical perspective so that we can design algorithms used in the wild reliably. Possible research directions include (i) understanding the vulnerabilities of other widely used algorithms, (ii) reducing the capabilities of the strategic actors, and (iii) using other approximation methods.
Appendix A Proof of Lemma III
Based on the definition (5), we have
$\|\overline{\pi}(z)-\overline{\pi}(\overline{z})$ $\|_{1}=\sum_{b\in B}|\overline{\pi}(z)(b)-\overline{\pi}(\overline{z})(b)|$
$=\sum_{b\in B}\left|\prod_{i\in I^{N}}\sigma(q^{i})(b^{i})-\prod_{i\in I^{N}}\sigma(\overline{q}^{i})(b^{i})\right|$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{b}\left(\prod_{i\neq\ell}\sigma(q^{i})(b^{i})\right)|\sigma(q^{\ell})(b^{\ell})-\sigma(\overline{q}^{\ell})(b^{\ell})|$
$\;\;\;\;+\sum_{b}\sigma(\overline{q}^{\ell})(b^{\ell})\left|\prod_{i\neq\ell}\sigma(q^{i})(b^{i})-\prod_{i\neq\ell}\sigma(\overline{q}^{i})(b^{i})\right|$
=$\stackrel{{\scriptstyle(b)}}{{=}}\|\sigma(q^{\ell})-\sigma(\overline{q}^{\ell})\|_{1}$
$\;\;\;\;+\sum_{b^{-\ell}}\left|\prod_{i\neq\ell}\sigma(q^{i})(b^{i})-\prod_{i\neq\ell}\sigma(\overline{q}^{i})(b^{i})\right|,$ (31)
where $(a)$ follows from the triangle inequality and by adding and subtracting $\left(\prod_{i\neq\ell}\sigma(q^{i})(b^{i})\right)\sigma(\overline{q}^{\ell})(b^{\ell})$ for some $\ell\in I^{N}$, and $(b)$ follows from the fact that the softmax, as described in (1) is a probability distribution. By following similar lines with (31), we can show that
$\|\overline{\pi}(z)-\overline{\pi}(\overline{z})\|_{1}$ $\leq\sum_{i\in I^{N}}\|\sigma(q^{i})-\sigma(\overline{q}^{i})\|_{1}$
$\leq\sum_{i}\sqrt{|B^{i}|}\cdot\|\sigma(q^{i})-\sigma(\overline{q}^{i})\|_{2}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{1}{\tau}\sum_{i}\sqrt{|B^{i}|}\cdot\|q^{i}-\overline{q}^{i}\|_{2}$
$\leq\frac{\sqrt{|B|}}{\tau}\sum_{i\in I^{N}}\|q^{i}-\overline{q}^{i}\|_{1}$
$=\frac{\sqrt{|B|}}{\tau}\cdot\|z-\overline{z}\|_{1},$ (32)
where $(a)$ follows from the $\frac{1}{\tau}$-Lipschitz continuity of the softmax function $\sigma(\cdot)$ with respect to $\|\cdot\|_{2}$, e.g., see [29, Proposition 4].
Appendix B Proof of Lemma III
By (4), the left-hand side of (16) can be written as
$|r^{j}(z,a)-r^{j}(\overline{z},a)|$ $=|u^{j}(a,\cdot)^{T}(\overline{\pi}(z)-\overline{\pi}(\overline{z}))|$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\|u^{j}(a,\cdot)\|_{\infty}\|\overline{\pi}(z)-\overline{\pi}(\overline{z})\|_{1}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}L_{0}\|z-\overline{z}\|_{1},$ (33)
where $(a)$ follows from the Hölder inequality, and $(b)$ follows from Lemma III and the definition of $L_{0}$.
Appendix C Proof of Lemma III
By the definition of the transition kernel $p(\cdot|\cdot)$, as described in Problem II, we can write the left-hand side of the inequality (18) as
$\Xi:=\left|\sum_{b}\overline{\pi}(z)(b)\cdot\eta(z_{+}^{ab})-\sum_{b}\overline{\pi}(\overline{z})(b)\cdot\eta(\overline{z}_{+}^{ab})\right|,$ (34)
where $z_{+}^{ab}=\{q_{+}^{i,ab}\}_{i\in I^{N}}$ is the next state given the current state and the action profile $(a,b)$. Correspondingly, $\overline{z}_{+}^{ab}=\{\overline{q}_{+}^{i,ab}\}_{i\in I^{N}}$ is the next state given the corresponding state $\overline{z}$ and the action profile $(a,b)$.
If we add and subtract $\sum_{b}\overline{\pi}(z)(b)\cdot\eta(\overline{z}_{+}^{ab})$, the triangle inequality yields that $\Xi\leq\Xi_{1}+\Xi_{2}$, where
$\Xi_{1}\coloneqq\left|\sum_{b}\overline{\pi}(z)(b)\cdot\left(\eta(z_{+}^{ab})-\eta(\overline{z}_{+}^{ab})\right)\right|,$ (35a)
$\Xi_{2}\coloneqq\left|\sum_{b}\Big{(}\overline{\pi}(z)(b)-\overline{\pi}(\overline{z})(b)\Big{)}\cdot\eta(\overline{z}_{+}^{ab})\right|.$ (35b)
Then, we can bound $\Xi_{1}$ by
$\Xi_{1}$ $\leq\sum_{b}\overline{\pi}(z)(b)\cdot\left|\eta(z_{+}^{ab})-\eta(\overline{z}_{+}^{ab})\right|$
$\stackrel{{\scriptstyle(a)}}{{\leq}}K\sum_{b}\overline{\pi}(z)(b)\cdot\|z_{+}^{ab}-\overline{z}_{+}^{ab}\|_{1}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}K\cdot\|z-\overline{z}\|_{1},$ (36)
where $(a)$ follows from the $K$-Lipschitz continuity of $\eta(\cdot)$ from the statement of the lemma, and $(b)$ follows from $\|\overline{\pi}(z)\|_{1}=1$ and the fact that
$\|z_{+}^{ab}-\overline{z}_{+}^{ab}\|_{1}=\sum_{i\in I^{N}}\sum_{b^{i}\in B^{i}}|q_{+}^{i,ab}(b^{i})-\overline{q}_{+}^{i,ab}(b^{i})|$
$=\sum_{i\in I^{N}}\left((1-\alpha)|q^{i}(b^{i})-\overline{q}^{i}(b^{i})|+\sum_{\tilde{b}^{i}\neq b^{i}}|q^{i}(\tilde{b}^{i})-\overline{q}^{i}(\tilde{b}^{i})|\right)$
$\leq\sum_{i\in I^{N}}\|q^{i}-\overline{q}^{i}\|_{1}.$ (37)
On the other hand, we can bound $\Xi_{2}$ by
$\Xi_{2}$ $\stackrel{{\scriptstyle(a)}}{{\leq}}\|\overline{\pi}(z)-\overline{\pi}(\overline{z})\|_{1}\cdot\max_{b\in B}|\eta(\overline{z}_{+}^{ab})|$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{\sqrt{|B|}}{\tau}M\cdot\|z-\overline{z}\|_{1}.$ (38)
where $(a)$ follows from the Hölder inequality and $(b)$ follows from Lemma III and the boundedness of $\eta(\cdot)$ from the statement of the lemma. Combining (36) and (38) lead to (18).